### v1

# The Peekaboo galaxy: new SALT spectroscopy and implications of archive HST data

Abstract
Context. The dwarf galaxy Peekaboo (HIPASSJ1131–31) was recently identified as a Local Volume (LV) gas-rich and ’eXtremely Metal-Poor’ (XMP) dIrr. Its gas metallicity is ZZ⊙/50, with 1 uncertainty range of [Z⊙/72–Z⊙/35]). Its the ”Tip of Red Giant Stars” (TRGB)-based distance is of 6.80.7 Mpc. The Hubble Space Telescope (HST) data for its individual stars reveal that its older RGB stars comprise a smaller part, while the majority of visible stars have ages of less than one to a few Gyr. Thus, Peekaboo dwarf can be considered as the nearest record-low-Z dwarf. As such, the galaxy deserves a deeper multi-method study, including properties of both, young massive stars and the fainter older population, and its ionised gas and the dominant baryonic component of Hi gas.
Aims. We aim to obtain the higher S-to-N SALT optical spectra of Peekaboo two H ii regions, in order to improve the accuracy of its gas O/H and to determine abundances of Ne, S, N, Ar. With the archive HST images, we aim to identify the hot massive stars (including exciting the two H ii regions) and the XMP supergiants as important targets for the follow-up studies of their evolution with upcoming Extremely Large Telescopes.
Methods. We use the direct (Te) method for the east H ii region, in which the [Oiii]4363Å line is well detected, to estimate its parameter 12+log(O/H). Since in the west H ii region the line [Oiii]4363Å is not detected, its O/H is estimated via the empirical ”strong-line” method of Izotov et al. The resulting value of O/H is very close to that in the east H ii region.
Results. The new spectroscopy of Peekaboo dwarf allows us to improve substantially the accuracy of its direct O/H estimate, which appears of 12+log(O/H) = 6.990.06 dex. The new data reveal that emission lines in the E region consist of two components with the velocity difference of 65 km s-1. The fainter, approaching, component can be related to a fast-moving WR star thrown from a cluster or a binary system. Using the HST magnitudes and colour , we identify tentative O-type and very hot candidate WO stars, which are likely the ionising stars of the studied H ii regions.
Conclusions. With the new optical spectra, the Pekaboo galaxy is confirmed as the lowest-metallicity dwarf in the Local Volume and the valuable object for indeep multi-method studies. Its most luminous stars are separated for the follow-up ground-based brightness monitoring and spectroscopy.
Key Words.:
galaxies: abundances – galaxies: individual: Peekaboo (HIPASSJ1131–31) – stars: massive – stars: evolution1 Introduction
The ’local’ gas-rich XMP (eXtremely Metal-Poor) dwarfs are interesting by themselves since they challenge the modern cosmological models of galaxy formation and evolution, allowing to check their predictions for the formation and survival of objects with the extreme properties. The gas-rich XMP dwarfs on the sum of their observational properties look like galaxies in the early stages of their evolution.
The existence of such unusual, very rare, so-called VYGs (Very Young Galaxies), was also predicted in N-body simulations (Tweed et al. 2018). VYGs are defined as galaxies formed the major part of their stellar mass within the last 1 Gyr. On the one hand, VYGs are expected to be very rare. On the other hand, their proportion in simulations is strongly depending on the type of Dark Matter (Cold DM versus Warm DM).
Besides the limitations on the cosmological models, which the statistics of XMP VYGs can imply, their study is of interest at least in two more areas.
Firstly, due to the expected similarity of their properties to those of young galaxies in the early Universe (very gas-rich and very metal-poor). Being at the distances of the LV and its environs, they allow one to get a much deeper insights in their properties and to tie them with the processes of galaxy formation in the early Universe.
The other important issue, related to stellar physics and galaxy evolution, is the basic understanding of formation and evolution of massive stars and their feedback into ISM in their parent galaxies (e.g. Eldridge, Stanway 2022). For cosmology of the early Universe, it is especially crucial to understand well properties and evolution of the massive stars with metallicities of tens - hundreds times lower than the solar one Z⊙. Such massive, very-low-Z stars can be found only in a handful XMP dwarf galaxies within the LV and its environs (Garcia et al. 2021). If identified, such XMP massive stars can be the primary targets for spectroscopic studies with the next generation ground-based telescopes. Meantime, the first steps in this direction are already presented for a nearby dIrr Leo A with Z 0.03 Z⊙ by Gull et al. (2022) and for four low-metallicity galaxies, including Leo P with Z 0.025 Z⊙ by Telford et al. (2024).
The Pekaboo galaxy, the optical counterpart of Hi source HIPASS J1131–31, got its name since it was hidden within the halo of a nearby bright star (Koribalski et al. 2018). The galaxy was recently described and studied by Karachentsev et al. (2023). Referring this paper, ”Peekaboo dwarf is a faint (MV = –11.27 mag), the Local Volume (LV) gas-rich dIrr at the TRGB-based distance of 6.8 Mpc (distance module = 29.2 mag). Its gas metallicity, derived with the direct Te method, appeared one of the lowest of the known so far (12+log(O/H) = 6.990.16 dex)”. The cited uncertainties of O/H correspond to the range of a factor of two: [Z⊙/72–Z⊙/35]). The authors also discuss the probable ”youth” of this galaxy, paying attention to the abnormally small population of aged RGB stars.
In this Letter, using the new SALT spectroscopy, we update the estimate of Peekaboo metallicity, increasing substantially its accuracy. Also, we use the archive HST images and the Extragalactic Distance Database (EDD)111https://edd.ifa.hawaii.edu photometry of its stars to identify the hot massive stars, ionising the two H ii regions, and several the brightest supergiants. The latter can be attractive targets for HST and ELT spectral study as the massive evolved stars with the record-low metallicities.
The layout of the Letter is as follows. In Sect. 2, the data and analysis are briefly described. Sect. 3 presents the main results. They are discussed in Sect. 4 and summarised in Sect. 5. The details of data and results are shown in Appendices. The adopted scale is 34 pc in 1 arcsec.
Date | Grating | Exposure | Spectral scale | Seeing | Spectral range | FWHM |
(sec) | (Å pixel-1) | (arcsec) | (Å) | (Å) | ||
2022 February 22 | PG900 | 13002 | 0.97 | 1.60 | 36506700 | 4.7 |
2022 February 22 | PG1800 | 13002 | 0.40 | 1.60 | 59507240 | 1.8 |
2022 February 26 | PG3000 | 11002 | 0.24 | 1.80 | 36904440 | 1.1 |
2022 February 27 | PG3000 | 11002 | 0.24 | 1.70 | 36904440 | 1.1 |
2022 March 09 | PG1800 | 11002 | 0.40 | 1.50 | 59507240 | 1.8 |
2024 March 06 | PG3000 | 11002 | 0.24 | 1.50 | 36904440 | 1.1 |
2024 March 07 | PG900 | 11002 | 0.97 | 1.40 | 36506700 | 4.7 |
Spectra for this date were originally obtained for paper Karachentsev et al. (2023). |
2 Observations and data analysis
2.1 SALT observations and analysis
We carried out long-slit optical spectroscopic observations of Peekaboo galaxy using the Robert Stobie Spectrograph (RSS; Burgh et al. 2003; Kobulnicky et al. 2003) installed at the Southern African Large Telescope (SALT; Buckley, Swart & Meiring 2006; O’Donoghue et al. 2006). Different VPH gratings were used with the long slit of 1.5″ by 8′ to cover the full spectral range from 3600 Å to 7250 Å. The position angle for all observations was 87.88° and slit position was exactly the same as it was shown in Figure 6 of Karachentsev et al. (2023), where reference star was used to put the slit in exactly the same position for different observations. All details are shown in Table 1, where all dates, seeing, covered spectral range, spectral scale and FWHM are shown. The final spectrum taken with grating PG0900 was used as a basic one, since it covers the most spectral range. Spectral data obtained using other gratings, were processed independently and emission lines were measured in the same spatial areas. After that, the line fluxes, measured on the spectra obtained with other gratings, were recalculated to the flux system obtained using the PG900 grating. Namely, the total fluxes of the bright lines, measured on spectra for both gratings, were used to determine the conversion factor, and then, the fluxes for the weaker lines were recalculated using this factor. Spectrophotometric standards were observed at twilight as part of the SALT standard calibration plan.
Given that SALT is equipped with an Atmospheric Dispersion Compensator (ADC), the effect of atmospheric dispersion at an arbitrary long-slit is negligible. Additionally, the SALT design entails the movement of the telescope’s pupil during the track and exposures, thereby resulting in a constant alteration of the telescope’s effective area. Therefore, accurate absolute photometry and spectrophotometry are not possible. However, the relative flux calibration could be used. Thus, the relative energy distribution in the spectra could be obtained with SALT data.
Fortunately, we have on the slit the reference star with the known magnitudes in several bands from the Legacy DR10 database Li et al. (2024), which we can use as a local standard. to derive the line fluxes and luminosity of H. See details in Appendix, Sect. B.
The primary data reduction was done with the SALT science pipeline (Crawford et al. 2010). The long-slit reduction was done with the RSS pipeline described by Kniazev (2022). The 1D spectra extraction, emission lines measurement and calculation of chemical abundances was done in the way described by Kniazev et al. (2008). The fully reduced 2D spectra of Peekaboo are shown Figure 2 with labelling of the most prominent lines. Since the RSS detector is a mosaic and consists of three CCDs, the positions of two gaps between these CCDs are also shown.
Errors in lines intensities have been propagated in the calculations of the reddening correction, electron temperatures and densities and have been propagated to the elemental abundance errors. The observed emission line intensities relatively to the F(H) line and the intensities corrected for interstellar extinction and underlying stellar absorption are presented in Table 2. The equivalent width (H) of the H line, the calculated absorption equivalent widths (abs) of the Balmer lines, the extinction coefficient (H) (a sum of the internal extinction in Peekaboo and foreground extinction in the Milky Way) and the subsequently recalculated value are also listed there.
2.2 Hubble Space Telescope data
HST data are well described by Karachentsev et al. (2023). Here we use the information on positions, magnitudes and colours of the hottest stars as the potential exciting stars of the observed H ii regions. The evolved luminous stars (supergiants) are also identified with this data. They should be the valuable material for a more advanced study of the lowest metallicity massive stars with the upcoming new facilities, including the ELT.
In Figure 1 we show two zoomed HST images. In the top panel, we show image in F606W filter, where all stars are marked with their numbers in Table 4. These is the same set of stars as in paper by Karachentsev et al. (2023). In the bottom panel we present the similarly zoomed colour HST image of Peekaboo, taken from the HST archive (Prog. SNAP 15992, PI R.B. Tully).
3 Results
3.1 Structure and kinematics of the main H ii regions
Analysis of the long-slit spectra revealed the presence of two distinct H ii regions. Figure 3 shows the profile of the H line along the slit, based on the spectral data obtained with the PG900 grating. The analysis clearly shows two regions, which we will refer to as regions E and W, since the slit was oriented roughly in the east-west direction across the Galaxy. These regions are prominent as two peaks in the H profile, although they are less pronounced in the continuum profile. Region W is fainter, with an extent of about 3 arcsec, while the brighter region E extends over about 5 arcsec. We defined approximate boundaries for these regions (as shown in the figure) in order to extract 1d spectra along the slit. The aim was to minimise spectral contamination between the regions while maximising the useful signal from the emission lines. The selected region boundaries are also shown in the Figure 2 as red horizontal lines and in the top panel of Figure 1 as red circles centred on the reference star with the corresponding radii. These boundaries were applied consistently to all the acquired spectral data. The one-dimensional spectra obtained by summing within these limits are shown in Figures 4 and 5.
A more detailed analysis of the extracted one-dimensional spectra uncovers an interesting peculiarity. Namely, all observed emission lines in the spectrum of region E appear double-component. Every line exhibits an asymmetric blue wing. To test the hypothesis that all lines represent two velocity components, we first fit each line with a single Gaussian, and then – with a two-Gaussian model. The related improvement of the reduced statistic is evaluated. The improvement is very significant. For various lines, it varies by a factor of 2 to 6.5. Examples of the fitting for H and [O iii]Å lines and H and H, for gratings PG900 and PG3000 are presented in Figures 6 and 7, respectively.
We calculate the radial velocities of two components in the E region and that of the W region, taking the weighted means on the three independent measurements with grisms PG900, PG1800 and PG3000. Each of these velocities, in turn, was derived from the redshifts of 6–7 the strongest emission lines. The final values of Vhel for these three subsystems are as follows: East(red): 72110 km s-1; East(blue): 6563 km s-1, West: 6976 km s-1. Thus, we accept for the further discussion these three Peekaboo ionised gas systems, of which East(red) and East(blue) are co-spatial within the limits of our modest spatial resolution.
3.2 New estimates of O/H and other elements
The measured and corrected line fluxes and the derived element abundances are presented in Tables 2 and 3. Here we briefly summarise the obtained results. In Sect. 3.1 we found that there are three separate subsystems on the slit.
For the East(red) subsystem, the line [O iii]4363 is detected at the level of 6.5. This allows us to directly determine electron temperature Te and derive Oxygen and the other element abundances via the direct method. We obtain its 12+log(O/H)(Te) = 6.990.06 dex. The two non-direct methods, described in Sect. D, give for the same subsystem the O/H values within 0.02–0.03 dex of the former. For the West and the East(blue) subsystems, the line [O iii]4363 is not detected. Their O/H are derived with the ’strong-line’ method of Izotov et al. (2019). For both objects, the derived O/H is within 0.02–0.03 dex from (O/H)(Te) of the East(red).
3.3 HST data on XMP stars in Peekaboo dwarf
Having in the EDD (Anand et al. 2021) the table with parameters of all 56 individual stars within the Peekaboo dwarf body, we sort them on the colour and separate the hottest stars with –0.25 (see Sect. 4.2), expected to ionise the surrounding gas. We then look for such stars which fall within (or near) the borders of the E and W H ii regions. They are the most likely stars, exciting these H ii regions.
There are also several supergiants in this XMP dwarf, which also are very interesting, since they are the nearest massive evolved stars with the record-low metallicity. They are attractive targets for the spectral study with the HST and the next generation ELTs. See top panel of Fig. 1 with the HST image of Peekaboo with positions of these stars indicated.
The majority of the rest stars with of 0.0–0.5 mag and M –2.5 mag, most probably belong to the Blue Helium Burning (BHeB) type with ages less than 200 Myr (McQuinn et al. 2012).
4 Discussion
4.1 Metallicity of Peekaboo in the context
Summarising the new determination of O/H in the Peekaboo dwarf, which appeared close to the earlier value, but with much improved accuracy, 12+log(O/H)dir = 6.99 0.06 dex, we compare this value to the other known dwarfs with the lowest metallicities in the LV. The best to our knowledge, within the LV (R11 Mpc), apart Peekaboo, there were found 10 XMP dwarfs with 12+log(O/H) 7.17 dex (or Z(gas) 0.03 Z⊙). Six of them are published in Pustilnik et al. (2010, 2021, 2024); Skillman et al. (2013); Izotov et al. (2012). See Table 6 in Appendix, where we compile their properties. Five of these six dwarfs reside within the nearby voids from Pustilnik et al. (2019). In the framework of the ongoing project ”Studying the void galaxies in the Local Volume”, we found 4 new LV XMP dwarfs, currently prepared for publication. Only four of the total 11 LV XMP galaxies (including Peekaboo) have their O/H derived with the direct method.
One dwarf, Leo P, with 12+log(O/H) =7.17 dex (Skillman et al. 2013) is the nearest known XMP galaxy, at 1.6 Mpc from the Local Group center. Its proximity allowed one to study its massive hot XMP star O7-8 V with the ground-based spectroscopy and to obtain the important constraints of its wind properties (Telford et al. 2023, 2024). The remaining Local Volume XMP dwarfs are situated much farther, at the distances of 7–11 Mpc. Thus, spectroscopy of their luminous stars should be available with HST and upcoming ELTs.
4.2 Hot and luminous stars. Blue outliers
The proximity of Peekaboo and the high resolution of the HST images allow one to identify the hottest massive stars, including stars ionising the E and W H ii regions. We separate in the diagram MV versus (Fig. 8), a region, where all O-stars (for luminosity classes from V to I) fall according to the state-of-art models at metallicity of Z = Z⊙/10, presented by Lorenzo et al. (2025). This is a rectangular with boundaries of from –0.32 to –0.25 mag, and MV between –3.30 and –6.15 mag. The list of these stars with their additional parameters is presented in Appendix. In Fig. 8, we also identify four supergiant stars, the evolved massive stars, conditionally with M –6.0 mag (or 21.4 23.5 mag). They probably represent the upper part of Blue Helium Burning (BHeB) stars according to McQuinn et al. (2012). They are interesting for detailed studies as the nearest evolved massive stars with the record-low metallicity.
In particular, they can be observed spectroscopically at the superb conditions (narrow slit and seeing of 0.5 arcsec) with the Very Large Telescope at ESO. In addition, there is a good opportunity to check their possible large-amplitude optical variability via monitoring of the entire galaxy and analysing light variations of the small regions around these supergiants. A successful example of monitoring of variability of the XMP LBV and other supergiants in H ii regions of the XMP galaxy DDO68 (12+log(O/H) = 6.98 – 7.3 dex, Pustilnik et al 2005; Izotov & Thuan 2007; Annibali et al. 2019) was recently presented in Pustilnik et al. (2024); Pustilnik, Perepelitsyna (2025).
Another type of evolved massive stars can show up in versus diagram as ’blue outliers’ (e.g. Lorenzo et al. 2022). These stars with M –3.5 to –5 mag and colours bluer than the blue edge for the Main Sequence stars (namely, of –0.32, Lorenzo et al. (2025)) can be the Helium-burning WO stars with the surface temperatures up to of 150–200 kK (Lorenzo et al. 2022, and references therein).
4.3 Properties of Peekaboo H ii regions
In Appendix section B we present the distribution of H flux along the slit, which looks like two close ’gaussians’ with FWHM of 3.4 and 3.7 arcsec. Their sizes along the slit (corrected for seeing of 1.5 arcsec) are FWHM 3.05 and 3.4 arcsec, or 116 pc and 128 pc. In the colour HST image one does not see such clear extended structures. These sizes probably represent the extent of the (smoothed due to seeing) collection of several smaller H ii regions, elongated along the galaxy major axis. Several of the brightest smaller H ii regions are marked in Fig. 1 and their parameters are summarised in Table 5.
As the two-component velocity profiles of all emission lines in the E H ii region evidence, this cell of ionised gas indeed has the complex origin. Roughly 30% of its nebular emission is related to the ionised gas, moving with the radial velocity of about –60 km s-1 (’blue’ component hereafter) relative to the main body, with the system velocity of V 716 km s-1 (Karachentsev et al. 2023). Both the radial velocities, of the E region more massive ’red’ velocity component, and that of the W H ii region are close to the system one.
This ionised gas should be connected with a fast-moving hot star. Such massive stars, products of co-evolution in binary systems are well known in the Galaxy. Due to the ’poor’ spatial resolution of the long-slit spectra, one can not directly address the size of the ’blue’ component. However, since the ’blue’ component gas moves together with its exciting massive star, one can suggest that this shell is related to the mass loss from the fast-moving star, occurred after the epoch when it was thrown out from a close binary (or a cluster). The abundance pattern in E(blue), similar to those in the other regions, indicates that its nebular emission originates in the swept-up ambient gas shell. The case of the complex nature of the E H ii region partly resembles the case of the H ii region A2 in IC4662. In this region, the echelle spectroscopy revealed three components, of which one, related to a known WR star, moves with the relative velocity of V 20 km s-1 (Kniazev 2025).
In Appendix (Sect. B) we estimate luminosities of the line H of the three emission-line subsystems. They are (7.5, 8.4 and 14.8)1036 erg s-1 for the W, E(blue) and E(red) subsystems, respectively. These luminosities can be used to estimate the conditional number of hot stars, ionising these regions. For further calculations, we adopt the relation between L(H) and the related flux of ionising quanta Q0: L(H) = 4.7610-13 Q0. Resulting values of Q0 for the three subsystem are as follows: (1.6, 1.8 and 3.2) 1049 phot s-1. Recently Telford et al. (2023) estimated the value of Q0 for the lowest accessible metallicity O7V star LP26 in Leo P (12+log(O/H)=7.17 dex). This number (3.740.67)1048 phot s-1, is 2.5 times lower than the value, usually adopted for O7V star with Z = Z⊙. That is, we need from 3 to 6 O7V XMP stars per a subsystem to provide the required amount of ionising photons.
The comparison of star positions (marked with their numbers) in the HST image (top panel of Fig. 1) and their positions in the diagram of Fig. 8, reveals within the borders of the E region only one certain luminous O star (No. 7) and about 1-mag fainter probable O-stars No. 5 and 6. On the other hand, two likely luminous WR stars (No. 2 and 4) reside in the middle of the E region. One more, candidate very hot WR star No. 1 is close to No. 2. In particular, their extremely blue colours are consistent with those of WO stars, for which their temperatures are estimated to be in the range of 150-210 kK (e.g. Tramper et al. 2015). Their parameter Q0 falls in the range of (0.8–3)1049 phot s-1. That is, even one such WO star is sufficient to provide ionising photons for each of the three H ii subsystems.
Since observations indicate that most of WR stars are situated on the edges of the related H ii regions (e.g., Crowther & Bibby 2009), this implies that they were ejected from the parent cluster. Therefore, one can expect that the ’fast-moving’ ’blue’ subsystem of E region is related with one of the above WR stars. The spectra of this region with a seeing of 0.5 arcsec could help to disentangle various subsystems within this region.
Apart these two elongated E and W complexes, numerous small blue non-stellar objects (’nebulosities’) are distributed along the whole dwarf body. Their typical sizes of 0.2–0.5 arcsec correspond to radii of 7–15 pc, or to Strömgren radii of 4–8 pc. Such small H ii regions are probably ionised and excited by individual hot massive stars. Their () colours are bluer than those of hot O-type stars. Such very blue () colours are predicted for the nebular emission in low-metallicity high-temperature H ii regions in Zackrisson et al. (2001); Anders & Fritze-v. Alvensleben (2003). In particular, Anders & Fritze-v. Alvensleben (2003) show that colour () for young starbursts with ZZ⊙/50 can vary in the range of –0.7 to –0.25 mag for ages from 5 to 15 Myr due to the very strong contribution of emission lines in -band. This is exactly the range which we see for the five regions N1–N5, shown in Fig. 8 and in Tab. 5. As the lowest metallicity individual H ii regions in the nearby Universe, they deserve spectroscopy with HST and/or ELTs.
5 Conclusions
Summarising presented above new data on the dwarf galaxy Peekaboo and their discussion, we arrive at the following conclusions:
1.
The new spectral data for the two H ii regions of Peekaboo dwarf allow us to uncover three ionised gas systems. Two systems, the E and W, are separated spatially, at 4 arcsec between their centres (140 pc), approximately along the major axis. In the E H ii region we see two ’co-spatial’ kinematically separated gas systems with the radial velocities of Vhel of 720 km s-1 and 654 km s-1(’red’ and ’blue’ components).
2.
Thanks to the improved S-to-N ratio in the principal faint line [Oiii]4363Å of the E region (red), we derive its O/H with the direct (Te) method: 12+log(O/H) = 6.990.06 dex. For the E(blue) subsystem and for the W H ii region, the line [Oiii]4363Å is undetected. Their O/H, derived via the empirical ”strong-line” method, appear very close to that of the E(red) subsystem with (O/H)(Te). Pekaboo galaxy is the lowest metallicity dwarf within the Local Volume and its environs, among galaxies with gas O/H derived via the direct method.
3.
We derive also abundances of the other heavy elements X (Ne, Ar, N, S), which are consistent with the average X/O, typical of galaxies in the lowest metallicity range.
4.
Using the EDD coordinates, magnitudes and colours of individual stars in the body of Peekaboo dwarf, we identify the hot massive stars (including candidate WO stars), tentatively ionising its H ii regions, as well as four supergiants in this galaxy. These supergiants, as the brightest and the lowest-metallicity evolved massive stars, are the valuable targets for the detailed studies with HST and the upcoming ELTs as well as for the ground-based monitoring of their possible large-amplitude variability. They can be good probes to confront them with the modern models of stellar evolution at Z 0.02 Z⊙.
Acknowledgements.
Spectral observations reported in this paper were obtained with the Southern African Large Telescope (SALT) programs 2021-1-MLT-001 and 2023-1-MLT-006. A. K. acknowledges support from the National Research Foundation (NRF) of South Africa. The study was conducted under the state assignment of Lomonosov Moscow State University. The work was also performed as part of the SAO RAS government contract by the Ministry of Science and Higher Education of the Russian Federation. We acknowledge comments and suggestions of the anonymous referee, which helped to improve the paper clarity. This work is based on observations made with the NASA/ESA HST on the program SNAP 15992, PI R.B. Tully. We acknowledge the use of the Extragalactic Distance Database and, in particular, its CMDs/TRGB part. This work has made use of data from the Legacy Survey (DR10). The Legacy Survey team is acknowledged for making their data products available to the scientific community.References
- Anand et al. (2021) Anand G.S., Rizzi L., Tully R.B., et al., 2021, AJ, 162, 80 (15pp)
- Anders & Fritze-v. Alvensleben (2003) Anders P. & Fritze-v. Alvensleben U., 2003, A&A, 401, 1063
- Annibali et al. (2019) Annibali F., La Torre V., Tosi M., et al., 2019, MNRAS, 482, 3892
- Buckley, Swart & Meiring (2006) Buckley, D.A.H., Swart, G.P., Meiring, J.G., 2006, SPIE, 6267
- Burgh et al. (2003) Burgh, E.B., Nordsieck, K.H., Kobulnicky, H.A., Williams, T.B., O’Donoghue, D., Smith, M.P., Percival, J.W., 2003, SPIE, 4841, 1463
- Crawford et al. (2010) Crawford S.M. et al., 2010, in Silva D. R., Peck A. B., Soifer B. T., Proc. SPIE Conf. Ser. Vol. 7737, Observatory Operations: Strategies, Processes, and Systems III. SPIE, Bellingham, p. 773725
- Crowther & Bibby (2009) Crowther P.A., Bibby J.L., 2009, A&A, 499, 455
- Eldridge, Stanway (2022) Eldridge J.J., Stanway E.E., 2022, ARAA, 60, 455
- Garcia et al. (2021) Garcia M., Evans C.J., Bestenlehner J.M. et al., 2021, Experimental Astronomy, 51:887-911
- Gull et al. (2022) Gull M., Weisz D.R., Senchyna P., et al., 2022, ApJ, 941, 206
- Izotov et al. (2006) Izotov Y.I., Stasińska G., Meynet G., et al., 2006, A&A, 448, 955
- Izotov & Thuan (2007) Izotov Y.I., Thuan T.X., 2007, ApJ, 665, 1115
- Izotov et al. (2007) Izotov Y.I., Thuan T.X., Guseva N.G., 2007, ApJ, 671, 1297
- Izotov et al. (2012) Izotov Y.I., Thuan T.X., Guseva N.G., 2012, A&A, 546, A122
- Izotov et al. (2019) Izotov Y.I., Guseva N.G., Frieke K.J., Henkel C., 2019, A&A, 523, A40
- Karachentsev et al. (2023) Karachentsev I.D., Makarova L.N., Koribalski B.S., Anand G.S., Tully R.B., Kniazev A.Y., 2023, MNRAS, 518, 5893
- Kniazev (2025) Kniazev A.Y., 2025, RAA, 25, id.045012, 16 pp.
- Kniazev et al. (2008) Kniazev A.Y., et al. 2008, MNRAS, 388, 1667
- Kniazev (2022) Kniazev A.Y., 2022, Astrophysical Bulletin, 77, 334-346
- Kobulnicky et al. (2003) Kobulnicky H.A., Nordsieck K.H., Burgh E.B., Smith M.P., Percival J.W., Williams T.B., O’Donoghue D., 2003, SPIE, 4841, 1634
- Koribalski et al. (2018) Koribalski B.S., Wang J., Kamphius P., et al. 2018, MNRAS, 478, 1611
- Li et al. (2024) Li C., Zhang Y., Cui C., Wei S., Zhang J., Zhao Y., Wu X.-B., et al., 2024, AJ, 168, 233
- Lorenzo et al. (2022) Lorenzo M., Garcia M., Najarro F., Herrero A., Cerviño M., Castro N., 2022, MNRAS, 516, 4164
- Lorenzo et al. (2025) Lorenzo M., Garcia M., Castro N., Najarro F., Cerviño M., Herrero A., Simón-Díaz S., 2025, MNRAS, 537, 1197
Lupton et al. (2005)
Lupton R., et al. 2005, https://www.sdss3.org/dr8/algorithms/
sdssUBVRITransform.php#Lupton2005 - Luridiana, Morisset, & Shaw (2015) Luridiana V., Morisset C., Shaw R. A., 2015, A&A, 573, A42
- McQuinn et al. (2012) McQuinn K.B.W., Skillman E.D., Dalcanton J.J., Cannon J.M., Dolphin A.E., Holtzman J., Weisz D.R., Williams B.F., 2012, ApJ, 759, id.77 (17pp)
- O’Donoghue et al. (2006) O’Donoghue D., et al. 2006, MNRAS, 372, 1510
- Pustilnik et al (2005) Pustilnik S., Kniazev A., Pramskij A., 2005, A&A, 443, 91
- Pustilnik et al. (2010) Pustilnik S.A., Tepliakova A.L., Kniazev A.Y., Burenkov A.N., 2010, MNRAS, 401, 333
- Pustilnik et al. (2019) Pustilnik S.A., Tepliakova A.L., Makarov D.I., 2019, MNRAS, 482, 4329
- Pustilnik et al. (2021) Pustilnik S.A., Egorova E.S., Kniazev A.Y., Perepelitsyna Y.A., Tepliakova A.L., Burenkov A.N., Oparin D.V., 2021, MNRAS, 507, 494
- Pustilnik et al. (2024) Pustilnik S.A., Kniazev A.Y., Tepliakova A.L., Perepelitsyna Y.A., Egorova E.S., 2024, MNRAS, 527, 11066
- Pustilnik et al. (2024) Pustilnik S.A., Perepelitsyna Y.A., Vinokurov A.S., et al. 2024, Astrophys.Bull., 79, 594 (arXiv:2411.07393)
- Pustilnik, Perepelitsyna (2025) Pustilnik S.A., & Perepelitsyna Y.A., 2025, A&A, 695, L7
- Schlafly, Finkbeiner (2011) Schlafly E.F., & Finkbeiner D.P., 2011, ApJ, 737, article id. 103
- Skillman et al. (2013) Skillman E., Salzer J.J., Berg D.A., et al., 2013, AJ, 146, 3
- Stasinska (2005) Stasinska G., 2005, A&A, 434, 507-520
- Telford et al. (2023) Telford O.G., McQuinn K.B.W., Chisholm J., Berg D.A., 2023, ApJ, 943:65 (17pp)
- Telford et al. (2024) Telford O.G., Chisholm J., Sander A.A.C., Ramachandran V., McQuinn K.B.W., Berg D.A., 2024, ApJ, 974:85 (22pp)
- Tramper et al. (2015) Tramper F., Straal S.M., Sanyal D., et al. 2015, A&A, 581, A110
- Tully et al. (2009) Tully R.B., et al., 2009, ApJ, 138, 323
- Tweed et al. (2018) Tweed D.P., Mamon G.A., Thuan T.X., Cattaneo A., Dekel A., Menci N., Calura F., Silk J., 2018, MNRAS, 477, 1427
- Zackrisson et al. (2001) Zackrisson E., Bergvall N., Olofsson K., Siebert A., 2001, A&A, 375, 814
Appendix A HST image of Peekaboo dwarf with the slit position
In the top panel of Fig. 1 we show the zoomed F606W HST image of Peekaboo dwarf. The position of the 1.5-arcsec-width long slit is superimposed. We also overplot approximate boundaries of W and E H ii regions with arcs, defined through the distance from the ’reference’ star. The hottest and supergiant stars are marked to show their location relative to W and E H ii regions.
The bottom panel presents the same scale colour HST image of the galaxy. The brightest stars are visible along with numerous blue non-stellar objects (’nebulosities’) with the sizes of a fraction of arcsec. The respective linear extent is 8 to 15 pc. These are likely small H ii regions, excited by individual massive hot stars. They are marked as N1–N7 in the top panel. Probably, the integrated nebular emission collected on the long slit, is just a sum of such smaller H ii regions, which, due to the smoothing effect of seeing, show up only as emission of regions with sizes of 100 pc (3 arcsec).
Appendix B 2D spectra, sizes of H ii regions and related issues
In Fig. 2 we present 2D spectra of Peekaboo with gratings PG900 and PG1800. Positions of E and W H ii regions along the slit as well as that of the reference star are indicated. The reference star (RA,Dec = 172.8887,–31.6742) has the available photometry in Legacy DR10 release (Li et al. 2024) with the following magnitudes: =19.93, =19.49, =19.35, =19.31. We use this photometry and its 1d spectrum (we extract its full extent along the slit and estimate the loss on the slit of 24%) to produce the flux-calibrated spectra of the E and W regions and to estimate fluxes in H lines. They, in turn, are used to estimate luminosities in this lines and the related parameters of exciting hot stars in these regions.
The fluxes of emission H line of W and E regions are 8.5 and 26.3, respectively (in units of 10-16 erg cm-2 s-1). The line flux for the E region comprises of two components, with the ’blue’ component part of 9.6 and for the ’red’ component of 16.7. The related line luminosities, after correction for the measured extinction in the H ii regions, C(H)0.18 dex, or factor of 1.6, are as follows: for W region: 7.51036, for E region: 23.21036 erg s-1, with division of (8.4 and 14.8)1036 for blue and red components, respectively.
Appendix C Extracted 1d spectra of E and W H ii regions
In Figure 4 we show 1d spectra of Peekaboo E and W regions with grism PG900, in which we mark the main emission lines used for the abundance determination (Hydrogen Balmer lines and lines of O, Ne, N, S, Ar ions). The spectra are flux-calibrated via the known Legacy DR10 photometry of the reference star. In Figure 5, a similar 1d spectrum is shown, with the higher spectral resolution with the use of grism PG3000. The faint auroral line [Oiii]4363 Å is well seen near H in the spectrum of the eastern H ii region.
Appendix D Line intensities and element abundances in the E and W H ii regions of Peekaboo dwarf
In Table 2 we present the measured relative line fluxes F() and corrected for extinction and underlying Balmer absorptions relative line intensities I() for the three subsystems described in Sect. 2.1 and 3.1. All details of line measurements and the subsequent determination of physical conditions and element abundances are described in Kniazev et al. (2008). The relevant collision strengths and transition probabilities were used from Stasinska (2005) and Izotov et al. (2006). The results of the abundance calculations, obtained using these programmes, were compared with those derived from the PyNeb package (Luridiana, Morisset, & Shaw 2015). The comparison demonstrates consistency within the stated uncertainties.
In the bottom of Table 2 we also present the derived extinction coefficient C(H), equivalent widths of Balmer absorptions EW(abs) in the underlying continuum and the equivalent widths of emission line H EW(H). These corrected line intensities I() are used for calculations of physical conditions and element abundances according to recipes from Izotov et al. (2006). The standard two-zone model is assumed with the internal hot zone, where oxygen is twice-ionised, and the outer colder zone, where oxygen is single-ionised. In Table 3, for the same three subsystems, we present Te for those two zones, and the estimate of electron density via the ratio of line fluxes in [Sii] doublet. Since for all three subsystems, this ratio exceeds the theoretical limit of 1.43, corresponding to the limit of the low density (1 cm-3), we adopt for all cases the conditional value of ne = 10 cm-3, typical of majority of extragalactic H ii regions. The uncertainties of Te are calculated according to Equation 3 from Kniazev et al. (2008) via the standard error transfer from the known uncertainties of the related fluxes of the Oxygen lines. The temperatures Te(O[ii]) and Te(S[iii]) are calculated using the approximation from Izotov et al. (2006).
The abundances of oxygen and other elements are derived with the direct method when the faint auroral line [O iii]4363 is available. When this line is below the noise level (as in the subsystems of E(blue) and Region W), we obtain the estimate of O/H with the modified semi-empirical method (mse) (suggested in Izotov & Thuan (2007) and modified with the account for effect of parameter O32, in Pustilnik et al. 2021). For the independent estimate of 12+log(O/H), we apply also the strong-line method of Izotov et al. (2019), which appears to demonstrate the best accuracy among all known empirical estimators. In particular, the independent check of its accuracy on the extended sample of the low-metallicity H ii regions in Pustilnik et al. (2021), resulted in its internal rms scatter of only 0.04 dex. As one can see, the estimates of O/H with the three different methods for these H ii regions give very close results.
The obtained abundance ratios of N/O, S/O, Ne/O, Ar/O in all three subsystems are consistent to each other. We can calculate their weighted averages as the characteristic values of the whole galaxy Peekaboo. The results are as follows: log(N/O)= –1.450.03, log(Ne/O)= –0.870.05; log(S/O) = -1.720.07, log(Ar/O) = –2.330.09. These averages are well consistent with the abundances ratios for the lowest metallicities, compiled by Izotov et al. (2006, in their Figure 10, left panel).
subsystem E(red) | subsystem E(blue) | Region W | ||||
(Å) Ion | F()/F(H) | I()/I(H) | F()/F(H) | I()/I(H) | F()/F(H) | I()/I(H) |
3727 [O ii] | 0.4280.023 | 0.4790.026 | 0.3940.039 | 0.4490.045 | 0.8210.034 | 0.9400.041 |
3868 [Ne iii] | 0.0880.009 | 0.0980.009 | 0.0810.017 | 0.0900.019 | 0.0620.013 | 0.0700.015 |
4101 H | 0.2520.010 | 0.2710.011 | 0.2530.017 | 0.2750.041 | 0.2840.012 | 0.3090.027 |
4340 H | 0.4270.014 | 0.4480.016 | 0.4580.025 | 0.4840.039 | 0.4690.017 | 0.4970.026 |
4363 [O iii] | 0.0320.005 | 0.0330.005 | — | — | — | — |
4861 H | 1.0000.027 | 1.0000.032 | 1.0000.047 | 1.0000.054 | 1.0000.027 | 1.0000.032 |
4959 [O iii] | 0.3300.022 | 0.3270.022 | 0.3700.037 | 0.3660.037 | 0.2310.020 | 0.2270.019 |
5007 [O iii] | 1.0430.043 | 1.0300.043 | 0.9710.060 | 0.9570.060 | 0.6320.031 | 0.6160.031 |
5876 He i | 0.0600.009 | 0.0560.009 | 0.0130.011 | 0.0120.010 | 0.0320.008 | 0.0280.007 |
6312 [S iii] | 0.0030.002 | 0.0030.002 | — | — | 0.0080.003 | 0.0070.003 |
6563 H | 3.0960.059 | 2.7400.058 | 3.1260.106 | 2.7140.101 | 3.1390.104 | 2.7140.098 |
6584 [N ii] | 0.0230.002 | 0.0200.002 | 0.0240.006 | 0.0210.005 | 0.0420.003 | 0.0360.003 |
6678 He i | 0.0260.003 | 0.0220.003 | 0.0290.006 | 0.0250.005 | 0.0190.003 | 0.0160.003 |
6717 [S ii] | 0.0460.004 | 0.0410.003 | 0.0370.006 | 0.0320.005 | 0.0700.005 | 0.0600.004 |
6731 [S ii] | 0.0210.003 | 0.0180.003 | 0.0260.006 | 0.0220.005 | 0.0380.003 | 0.0330.003 |
7065 He i | 0.0250.003 | 0.0220.002 | — | — | 0.0190.003 | 0.0160.003 |
7136 [Ar iii] | 0.0200.003 | 0.0170.003 | 0.0090.004 | 0.0070.003 | 0.0110.003 | 0.0090.003 |
C(H) dex | 0.160.03 | 0.180.04 | 0.190.04 | |||
EW(abs) Å | 0.000.29 | 0.000.26 | 0.000.33 | |||
EW(H) Å | 18 2 | 12 3 | 193 |
Value | subsystem E(red)(4363) | subsystem E(blue) | Region W |
PG900+PG3000+PG1800 | PG900+PG3000+PG1800 | PG900+PG3000+PG1800 | |
(OIII)(K) | 19,6821990 | 22,0322316 | 22,4882285 |
(OII)(K) | 15,582208 | 16,209508 | 16,305463 |
(SIII)(K) | 18,5421327 | 19,8771087 | 20,082985 |
(SII)(cm-3) | 1010 | 1010 | 1010 |
O+/H+(105) | 0.3870.026 | 0.3230.044 | 0.6640.062 |
O++/H+(105) | 0.6010.128 | 0.4680.098 | 0.2890.057 |
O/H(105) | 0.9880.131 | 0.7900.107 | 0.9530.084 |
12+log(O/H) (dir) | 6.990.06 | … | … |
12+log(O/H)(mse,c) | 6.960.11 | 6.940.11 | 7.020.10 |
12+log(O/H)(s) | 7.010.04 | 6.980.05 | 7.020.05 |
N+/H+(107) | 1.4290.121 | 1.3430.283 | 2.3040.202 |
ICF(N) | 2.573 | 2.471 | 1.368 |
N/H(105) | 0.040.00 | 0.030.01 | 0.030.00 |
12+log(N/H) | 5.570.04 | 5.520.09 | 5.500.04 |
log(N/O) | -1.430.07 | -1.380.11 | -1.480.05 |
Ne++/H+(105) | 0.1280.033 | 0.0920.028 | 0.0680.020 |
ICF(Ne) | 1.167 | 1.174 | 1.321 |
Ne/H(105) | 0.150.04 | 0.110.03 | 0.090.03 |
12+log(Ne/H) | 6.170.11 | 6.030.13 | 5.960.13 |
log(Ne/O) | -0.820.12 | -0.860.15 | -1.020.14 |
S+/H+(107) | 0.5290.040 | — | 0.7720.057 |
S++/H+(107) | 0.8460.553 | — | 1.6200.679 |
ICF(S) | 0.970 | — | 0.826 |
S/H(107) | 1.330.54 | — | 1.980.56 |
12+log(S/H) | 5.120.18 | — | 5.300.12 |
log(S/O) | -1.870.18 | — | -1.680.13 |
Ar++/H+(107) | 0.4890.092 | 0.1960.080 | 0.2340.069 |
ICF(Ar) | 1.075 | 1.074 | 1.103 |
Ar/H(107) | 0.530.10 | 0.210.09 | 0.260.08 |
12+log(Ar/H) | 4.720.08 | 4.320.18 | 4.410.13 |
log(Ar/O) | -2.270.10 | -2.580.19 | -2.570.13 |
Abundances of O, N, Ne, and Ar for subsystems E(blue) and Region W, and of S for Region W are calculated with Te, | |||
adopted in the modified semi-empirical method (mse) (Pustilnik et al. 2021). Relative values X/O are direct results of calculations. | |||
12+log(O/H)(mse) is, however, corrected upward by 0.04 dex for consistency with O/H(Te) zero point (see Pustilnik et al. 2021). |
Appendix E HST data: identification of the XMP massive hot stars and supergiants
We plot in Fig. 8 positions of all HST stars within the Peekaboo body in the diagram MV versus colour . We mark by their numbers from Table 4 all ’blue’ hot stars and the four supergiants with M 6.0 mag. Here, for both MV and , only extinction in the Galaxy is accounted for, corresponding to E(B-V) = 0.062 (Schlafly, Finkbeiner 2011), and the related AV = 0.186 mag and E() = 0.066 mag. As one can see in the bottom of Table 2, C(H) for all three systems is the same within rather small uncertainties. We adopt its average value of 0.17. Then, for stars residing within the borders of these H ii regions, one can expect the elevated extinction and reddening, in addition to the adopted Milky Way values, AV = 0.16 mag and E()=0.057 mag. In fact, we do not know how this extinction is distributed within the E and W regions. If it is localised within the ’blue nebulosities’, then this ’reddening’ correction is probably applicable only to stars residing close to these ’blue nebulosities’.
Therefore, looking at the marked positions of blue and bluish stars in the Peekaboo dwarf image in Fig. 1 (top), one can understand where these stars shift in Fig. 8 with the account of the additional extinction (its value is shown by the ”arrow” in the plot). Curiously, but no stellar objects are identified within the W region. At the same time, two very blue and rather luminous ’nebulosities’, N4 and N5 can be the main contributors to the visible nebular emission of the W H ii region. As for the E H ii region, there are several stars, for which the additional extinction can take place. These are stars No. 1, 2, 4, 5, 6, 7, 8. Stars No. 10, 15, 18 and 20 reside in the ’narrow’ layer between the E and W H ii regions and also can be affected by the additional extinction. So, for some of the discussed blue stars, their colour index can be bluer than that shown in Fig. 8. This implies that some of visible hot stars within the E region can shift after the colour correction to the area of ’blue outliers’, where the very hot helium burning stars similar to WO stars reside (Lorenzo et al. 2022).
id | MV,0 | x | y | ||||||
mag | mag | mag | mag | mag | mag | mag | px | px | |
01 | -0.819 | 25.701 | 0.060 | 0.206 | -3.685 | -0.885 | 0.291 | 1945.99 | 2406.69 |
02 | -0.541 | 25.009 | 0.038 | 0.096 | -4.377 | -0.607 | 0.136 | 1938.69 | 2411.02 |
03 | -0.501 | 25.981 | 0.074 | 0.202 | -3.405 | -0.567 | 0.286 | 1908.78 | 2438.43 |
04 | -0.302 | 25.142 | 0.039 | 0.089 | -4.244 | -0.368 | 0.126 | 1926.85 | 2396.91 |
05 | -0.284 | 26.080 | 0.078 | 0.186 | -3.306 | -0.350 | 0.263 | 1922.01 | 2382.68 |
06 | -0.262 | 26.003 | 0.072 | 0.166 | -3.383 | -0.328 | 0.235 | 1919.79 | 2394.47 |
07 | -0.211 | 24.999 | 0.036 | 0.075 | -4.387 | -0.277 | 0.106 | 1954.70 | 2435.43 |
08 | -0.174 | 23.030 | 0.012 | 0.020 | -6.356 | -0.240 | 0.028 | 1932.66 | 2406.53 |
09 | -0.133 | 26.402 | 0.098 | 0.204 | -2.984 | -0.199 | 0.288 | 1979.08 | 2378.20 |
10 | -0.127 | 24.990 | 0.034 | 0.067 | -4.396 | -0.193 | 0.095 | 1900.32 | 2388.97 |
11 | -0.106 | 26.500 | 0.101 | 0.218 | -2.886 | -0.172 | 0.308 | 1936.33 | 2336.98 |
12 | -0.097 | 24.163 | 0.021 | 0.039 | -5.223 | -0.163 | 0.055 | 1907.77 | 2413.47 |
13 | -0.090 | 25.576 | 0.052 | 0.099 | -3.810 | -0.156 | 0.140 | 1906.00 | 2448.44 |
14 | -0.079 | 26.130 | 0.077 | 0.152 | -3.256 | -0.145 | 0.215 | 1903.38 | 2446.73 |
15 | -0.070 | 25.619 | 0.051 | 0.101 | -3.767 | -0.136 | 0.143 | 1903.17 | 2373.19 |
16 | -0.063 | 25.522 | 0.051 | 0.098 | -3.864 | -0.129 | 0.139 | 1900.36 | 2407.79 |
17 | -0.035 | 25.946 | 0.066 | 0.126 | -3.440 | -0.101 | 0.178 | 1897.10 | 2412.65 |
18 | 0.008 | 23.265 | 0.012 | 0.020 | -6.121 | -0.058 | 0.028 | 1917.10 | 2370.44 |
19 | 0.024 | 26.408 | 0.094 | 0.172 | -2.978 | -0.042 | 0.243 | 1899.36 | 2425.37 |
20 | 0.030 | 22.703 | 0.009 | 0.014 | -6.683 | -0.036 | 0.020 | 1925.07 | 2369.57 |
21 | 0.053 | 25.430 | 0.048 | 0.084 | -3.956 | -0.013 | 0.119 | 1951.46 | 2410.61 |
22 | 0.064 | 26.254 | 0.092 | 0.155 | -3.132 | -0.002 | 0.219 | 1941.83 | 2406.24 |
23 | 0.078 | 25.810 | 0.057 | 0.020 | -3.576 | 0.012 | 0.028 | 1898.75 | 2458.86 |
24 | 0.106 | 25.133 | 0.037 | 0.060 | -4.253 | 0.040 | 0.085 | 1921.18 | 2411.16 |
25 | 0.112 | 23.998 | 0.018 | 0.029 | -5.388 | 0.046 | 0.041 | 1909.35 | 2342.51 |
26 | 0.160 | 25.946 | 0.071 | 0.113 | -3.440 | 0.094 | 0.160 | 1953.89 | 2439.19 |
27 | 0.169 | 21.406 | 0.004 | 0.006 | -7.980 | 0.103 | 0.008 | 1942.40 | 2466.17 |
28 | 0.177 | 26.496 | 0.095 | 0.157 | -2.890 | 0.111 | 0.222 | 2011.78 | 2403.17 |
29 | 0.201 | 26.339 | 0.086 | 0.139 | -3.047 | 0.135 | 0.197 | 1940.86 | 2340.86 |
30 | 0.202 | 25.127 | 0.039 | 0.059 | -4.259 | 0.136 | 0.083 | 1911.97 | 2413.27 |
31 | 0.209 | 26.731 | 0.117 | 0.188 | -2.655 | 0.143 | 0.266 | 1921.02 | 2344.43 |
32 | 0.243 | 26.328 | 0.083 | 0.183 | -3.058 | 0.177 | 0.259 | 1875.11 | 2401.46 |
33 | 0.243 | 26.808 | 0.124 | 0.204 | -2.578 | 0.177 | 0.288 | 1896.78 | 2302.76 |
34 | 0.255 | 24.819 | 0.028 | 0.042 | -4.567 | 0.189 | 0.059 | 1893.44 | 2308.65 |
35 | 0.304 | 26.662 | 0.108 | 0.161 | -2.724 | 0.238 | 0.228 | 1904.78 | 2340.19 |
36 | 0.389 | 26.912 | 0.131 | 0.185 | -2.474 | 0.323 | 0.262 | 1941.23 | 2345.01 |
37 | 0.414 | 26.548 | 0.098 | 0.134 | -2.838 | 0.348 | 0.189 | 1923.55 | 2384.49 |
38 | 0.490 | 26.842 | 0.115 | 0.151 | -2.544 | 0.424 | 0.214 | 1878.62 | 2421.48 |
39 | 0.527 | 25.128 | 0.034 | 0.041 | -4.258 | 0.461 | 0.058 | 1913.15 | 2427.36 |
40 | 0.567 | 26.332 | 0.081 | 0.097 | -3.054 | 0.501 | 0.137 | 1905.14 | 2404.46 |
41 | 0.653 | 26.744 | 0.100 | 0.118 | -2.642 | 0.587 | 0.167 | 1900.86 | 2295.04 |
42 | 0.686 | 26.497 | 0.086 | 0.099 | -2.889 | 0.620 | 0.140 | 1933.49 | 2343.37 |
43 | 0.722 | 27.444 | 0.178 | 0.192 | -1.942 | 0.656 | 0.272 | 1882.12 | 2400.30 |
44 | 0.729 | 26.966 | 0.129 | 0.133 | -2.420 | 0.663 | 0.188 | 1925.61 | 2392.75 |
45 | 0.748 | 26.776 | 0.098 | 0.106 | -2.610 | 0.682 | 0.150 | 1982.51 | 2436.69 |
46 | 0.787 | 26.495 | 0.082 | 0.084 | -2.891 | 0.721 | 0.119 | 1873.15 | 2391.30 |
47 | 0.790 | 27.057 | 0.127 | 0.128 | -2.329 | 0.724 | 0.181 | 1868.12 | 2413.24 |
48 | 0.806 | 26.608 | 0.088 | 0.093 | -2.778 | 0.740 | 0.132 | 1977.93 | 2313.70 |
49 | 0.900 | 27.425 | 0.153 | 0.149 | -1.961 | 0.834 | 0.211 | 1854.12 | 2307.70 |
50 | 0.945 | 24.232 | 0.017 | 0.016 | -5.154 | 0.879 | 0.023 | 1936.36 | 2421.46 |
51 | 0.953 | 27.248 | 0.134 | 0.124 | -2.138 | 0.887 | 0.175 | 1869.25 | 2406.83 |
52 | 1.028 | 27.403 | 0.158 | 0.139 | -1.983 | 0.962 | 0.197 | 1988.41 | 2337.56 |
53 | 1.054 | 26.688 | 0.082 | 0.074 | -2.698 | 0.988 | 0.105 | 1853.88 | 2431.71 |
54 | 1.239 | 27.350 | 0.140 | 0.101 | -2.036 | 1.173 | 0.143 | 1865.83 | 2343.97 |
55 | 1.310 | 24.629 | 0.020 | 0.016 | -4.757 | 1.244 | 0.023 | 1931.53 | 2330.40 |
56 | 1.483 | 25.406 | 0.033 | 0.022 | -3.980 | 1.417 | 0.031 | 1920.05 | 2350.67 |
id | MV,0 | RA | DEC | ||||||
mag | mag | mag | mag | mag | mag | mag | degrees | degrees | |
N 1 | -0.35 | 23.89 | 0.02 | 0.03 | -5.496 | -0.416 | 0.04 | 172.8952917 | -31.6739733 |
N 2 | -0.24 | 24.17 | 0.04 | 0.05 | -5.216 | -0.306 | 0.06 | 172.8947500 | -31.6739306 |
N 3 | -0.61 | 23.80 | 0.02 | 0.03 | -5.586 | -0.676 | 0.04 | 172.8952458 | -31.6744069 |
N 4 | -0.40 | 24.03 | 0.02 | 0.04 | -5.356 | -0.466 | 0.04 | 172.8940750 | -31.6740158 |
N 5 | -0.32 | 24.37 | 0.03 | 0.04 | -5.016 | -0.386 | 0.05 | 172.8936750 | -31.6741089 |
N 6 | -0.25 | 24.53 | 0.04 | 0.06 | -4.856 | -0.316 | 0.07 | 172.8950989 | -31.6745340 |
N 7 | -0.36 | 24.48 | 0.03 | 0.04 | -4.906 | -0.426 | 0.05 | 172.8945400 | -31.6746452 |
Name | Alter.name | 12+log(O/H) | Method | Vhel | MB,0 | Coord. J2000 | Ref. | |
dex | km s-1 | Mpc | mag | of (O/H) | ||||
HIPASSJ1131-31 | Peekaboo | 6.990.06 | Te | 716 | 6.8 | –11.25 | J113134.6-314026 | 1 |
AGC227973 | 7.070.04 | s.l. | 675 | 9.1 | –10.42 | J125039.9+052052 | 2 | |
SDSSJ0926+3343 | AGC194054 | 7.120.02 | Te | 532 | 10.8 | –12.92 | J092610.1+334312 | 3 |
SDSSJ1444+4242 | 7.140.04 | s.l. | 634 | 10.9 | –10.54 | J144449.8+424254 | 4 | |
PGC000083 | AGC103567 | 7.150.03 | s.l. | 542 | 9.1 | –12.89 | J000106.5+322241 | 4 |
SDSSJ1056+3608 | AGC205685 | 7.160.07 | Te | 572 | 9.2 | –11.60 | J105640.4+360828 | 5 |
HIJ1021+1805 | Leo P | 7.170.04 | Te | 264 | 1.6 | –09.11 | J102144.8+180520 | 6 |
1. This paper; 2. Pustilnik et al. (2024); 3. Pustilnik et al. (2010); | ||||||||
4. Pustilnik et al. (2021); 5. Izotov et al. (2012); 6. Skillman et al. (2013) |